OpenStudy (anonymous):

solve the system of equations x+3y=4 x=7-3y

OpenStudy (katrinakaif):

plug it in the equation so: 7-3y) + 3y = 4 Doyou know hwo to solve the rest?

OpenStudy (anonymous):

I think it is y=11 over 6 is that correct?

OpenStudy (katrinakaif):

Bring the 7 to the other side by substracting, so: -3y +3y = -3 Then combine lke terms 0y = -3 Final y = -3 Plug it into the first equation or second equation. x = 7-3y x = 7 - 3(-3) x = 7 + 9 x = 16

OpenStudy (anonymous):

The system is inconsistent.

OpenStudy (anonymous):

x+3y = 4 x+3y = 7 is inconsistent

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

There are no solutions.

OpenStudy (katrinakaif):

You have to use substituion in this alchemista

OpenStudy (anonymous):

Alchemista is right

OpenStudy (katrinakaif):

Ur welcome Babyfroggy :)

OpenStudy (anonymous):

Sorry but there are no solutions.

OpenStudy (katrinakaif):

Babyfroggy you got to plug it ack in the solution and see if it works.. I haven't done that yet

OpenStudy (anonymous):

If you want to think about it geometrically, it would be the same as saying that the same line intersects. There is no intersection, there is no solution.

OpenStudy (anonymous):

You need to rearrange the system in its proper form to see this x+3y=4 x=7-3y becomes x+3y = 4 x+3y = 7

OpenStudy (katrinakaif):

Sorry i forgot to go in and check it back again. The answer would be no solution if it does not work.

OpenStudy (anonymous):

Thanks everyone